The number of computer malfunctions per day is recorded for 260 days with the

following results.


Number of malfunctions (x i ) 0 1 2 3 4 5


Number of days(f i ) 77 90 55 30 5 3


Test the goodness of fit of an appropriate probability model.


1
Expert's answer
2021-02-03T03:55:05-0500

An appropriate probability model could be a Poisson distribution on the basis of following assumptions:

malfunctions occur independently,

simultaneous malfunctions are impossible,

malfunctions occur randomly in time,

malfunctions occur uniformly (mean number for time period

proportional to the period length).


A Poisson distribution has one parameter, "\lambda," which is the mean (and also the variance).


"mean=\bar{x}=\dfrac{\sum f_ix_i}{\sum f_i}"

"\sum f_ix_i=77(0)+90(1)+55(2)+30(3)+5(4)"


"+3(5)=325"




"\sum f_i=77+90+55+30+5+3=260"


"mean=\bar{x}=\dfrac{325}{260}=1.25"

"\sum f_ix_i^2=77(0)^2+90(1)^2+55(2)^2+30(3)^2"


"+5(4)^2+3(5)^2=735"

"Var(X)=s^2=\dfrac{735}{260}-1.25^2=1.264423"

Hence with "\lambda=1.25"


"P(X=x)=\dfrac{e^{-1.25}(1.25)^x}{x!}, x=0,1,2,3,..."

which gives the following probabilities and expected frequencies

(260"\times" probability)


"\begin{matrix}
 x_i & P(X=x_i) & E_i \\ \hline \\
 0 & 0.2865 & 74.5 \\
 1 & 0.3581 & 93.1 \\
 2 & 0.2238 & 58.2 \\
 3 & 0.0933 & 24.2 \\
 4 & 0.0291 & 7.6 \\
 5 & 0.0073 & 1.9 \\
 \geq6 & 0.0019 & 0.5 \\
& 1.0000 & 260.0
\end{matrix}"

Since a Poisson distribution is valid for all positive values of "X," the additional class "X\geq6" is necessary, and

"P(X\geq6)=1-P(X\leq6)"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c: c}
 x_i & O_i=f_i & P(X=x_i) & E_i \\ \hline
 0 & 77 & 0.2865 & 74.5 \\
 \hdashline
 1 & 90 & 0.3581 & 93.1 \\
 \hdashline 
2 & 55 & 0.2235 & 58.2 \\
 \hdashline
3 & 30 & 0.0933 & 24.2 \\
 \hdashline
\geq4 & 8 & 0.0383 & 10.0 \\
\hdashline
& & 1.0000 & \\
\end{array}"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c: c}
 x_i & O_i-E_i & (O_i-E_i)^2 & \dfrac{(O_i-E_i)^2}{E_i} \\ \hline
 0 & 2.5 &6.25 & 0.084 \\
 \hdashline
 1 & -3.1 & 9.61 & 0.103 \\
 \hdashline 
2 & -3.2 & 10.24 & 0.176 \\
 \hdashline
3 & 5.8 & 33.64 & 1.390 \\
 \hdashline
\geq4 & -2.0 & 4.00 & 0.400 \\
\hdashline
& & & 2.153\\
\end{array}"

"H_0:" number of daily malfunctions is "\sim" Poisson

"H_1:" number of daily malfunctions is not"\sim" Poisson


Significance level, "\alpha=0.05"

Degrees of freedom, "\nu=5-2=3"

5 classes: "0,1,2,3,\geq4"

2 constrants: "\sum E_i=\sum O_i" and "\sum E_ix_i=\sum O_ix_i"

from estimation of "\lambda"

Critical region "\chi^2>7.815"


Test statistic is


"\chi^2=\sum\dfrac{(O_i-E_i)^2}{E_i}=2.153"

This value does not lie in the critical region.

There is no evidence, at the 5% significant level, to suggest that the number

of computed malfunctions per day does not have a Poisson distribution.